OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and third equations, then the second and third equations. x – y + 2z = –2 2x + 2y + z = 7 3x + 3y – z = 3 last one..thnx

OpenStudy (anonymous):

can someone plzz help me?? HERO??

OpenStudy (anonymous):

mult [3] by 2 and add to [1] x – y + 2z = –2 [1] 6x + 6y –2z = 6 [3]'=[3]*2 =========== 7x+5y=4 [1]' Add the given [3] to [2] 2x + 2y + z = 7 [2] 3x + 3y – z = 3 [3] ========== 5x+5y=10 [2]' The new system is the 2x2 system 7x+5y=4 [1]' 5x+5y=10 [2]' I checked this one wolfram and the answers match up

OpenStudy (anonymous):

THANK YOU SOO MUCH FOR YOUR HELP ON BOTH OF MY QUESTIONS =)